OpenStudy (anonymous):

13y-2/4y - 13y+14/4y

OpenStudy (anonymous):

-4/y

OpenStudy (anonymous):

@neel how did u get 12 ? O.o

OpenStudy (anonymous):

m getin 3/y

OpenStudy (anonymous):

i took whole upon 4y

OpenStudy (anonymous):

\[(13y-2)/4y - (13y+14)/4y = (-2-14)/4y=-16/4y=-4/y\]

OpenStudy (anonymous):

ok... was bit wrong in inerruption

OpenStudy (anonymous):

ur rite